### v1

@@ 3.3 The Signal Decay Bound
In applications: $H_{0}$ corresponds to “step $t$ is valid” (the transition $K_{t-1}^{(0)}$ produces correct states), while $H_{1}$ corresponds to “step $t$ contains an error” (the transition $K_{t-1}^{(1)}$ may produce erroneous states). In manufacturing: $H_{0}$ means stage $t$ operated correctly; $H_{1}$ means a defect was introduced at stage $t$. The shared-suffix condition isolates step $t$ as the sole source of difference: downstream stages process whatever they receive in the same way under both hypotheses, so any difference in the outcome distribution is attributable entirely to step $t$. Errors introduced at step $t$ may be corrected or may cascade through subsequent stages, but the stages themselves operate identically. We index step $t$ as the abstract state $Z_{t}$ immediately after the $t$-th transition, so HtH-t counts the number of transitions remaining to reach the outcome.
@@
$\mathrm{err}^{*}(P^{(0)},P^{(1)}):=\inf_{\psi}\left(\mathbb{P}_{P^{(0)}}(\psi=1)+\mathbb{P}_{P^{(1)}}(\psi=0)\right)=1-d_{\mathrm{TV}}(P^{(0)},P^{(1)}),$ (11)
where the infimum is over all measurable tests $\psi:\mathcal{Y}\to\{0,1\}$ and $\psi=i$ denotes “accept $H_{i}$.” The hypotheses are $(n,\epsilon)$-distinguishable if the total testing error using $n$ i.i.d. samples is at most $\epsilon$.
Thus achieving total error at most $\epsilon$ requires $d_{\mathrm{TV}}(P^{(0)},P^{(1)})\geq 1-\epsilon$.
@@
    Sample complexity lower bound: For any target error $\epsilon\in(0,1/2)$, in the regime $\eta^{H-t}\Delta^{2}\leq 1$, any test achieving minimax error at most $\epsilon$ requires
    $n\geq\frac{(1-\epsilon)^{2}}{\eta^{H-t}\cdot\Delta^{2}}$ (13)
    i.i.d. outcome samples. When $\eta^{H-t}\Delta^{2}>1$, this lower bound becomes $O(1)$; in this regime, the outcome distributions are well-separated and $O(1)$ samples suffice to achieve any constant error rate.
@@
    $n=\Omega\left(\left(\frac{1}{\eta}\right)^{H-t}\right).$ (14)
@@
For initial divergence $\Delta^{2}$, sample budget $n$, and any target error $\epsilon\in(0,1/2)$, a necessary condition for achieving minimax testing error at most $\epsilon$ is that the step lies within the $\epsilon$-critical horizon:
$H-t\leq H_{\mathrm{crit}}(\epsilon):=\frac{\ln\!\bigl(n\Delta^{2}/(1-\epsilon)^{2}\bigr)}{\ln(1/\eta)}=\frac{\ln(n\Delta^{2})-2\ln(1-\epsilon)}{\ln(1/\eta)}.$ (15)
Equivalently, steps more than $H_{\mathrm{crit}}(\epsilon)$ away from the outcome cannot be reliably evaluated from outcome labels alone. For intuition, $H_{\mathrm{crit}}\approx\ln(n\Delta^{2})/\ln(1/\eta)$ when $\epsilon$ is a fixed constant.
@@
Consider contraction coefficient $\eta=0.9$, initial divergence $\Delta^{2}=0.1$, and sample budget $n=10^{6}$. The critical horizon is:
$H_{\mathrm{crit}}=\frac{\ln(10^{6}\times 0.1)}{\ln(1/0.9)}=\frac{\ln(10^{5})}{\ln(1.111)}\approx\frac{11.5}{0.105}\approx 110.$ (16)
With one million outcome samples, we can reliably evaluate steps up to 110 steps before the outcome. Steps further back are beyond the reach of any test.
With stronger contraction $\eta=0.7$ (more diverse sampling):
$H_{\mathrm{crit}}=\frac{11.5}{\ln(1/0.7)}\approx\frac{11.5}{0.357}\approx 32.$ (17)
Stronger mixing dramatically reduces the horizon over which credit assignment is possible. The same sample budget now reaches only 32 steps back. The formula depends logarithmically on $n$ and $\Delta^{2}$ but much more strongly on $\eta$: doubling the sample budget extends the horizon by $\ln 2/\ln(1/\eta)$ steps (about $6.6$ when $\eta=0.9$), while reducing $\eta$ from $0.9$ to $0.7$ cuts it by a factor of three.
Is the exponential rate tight? The following proposition constructs a Bernoulli instance where the likelihood ratio test matches the lower bound up to constants.
@@
Then the $\chi^{2}$-divergence between outcome distributions is $\chi^{2}(P^{(0)}_{R}\|P^{(1)}_{R})=\Theta(\eta^{H-t}\Delta^{2})$, and the standard likelihood ratio test achieves minimax error $\epsilon$ with sample complexity $n=O(1/(\eta^{H-t}\Delta^{2}))$.
Together, Theorem 3.9 and Proposition 3.12 show that the exponential rate $(1/\eta)^{H-t}$ is both necessary and sufficient: the Signal Decay Bound is tight up to constants. The exponential barrier is therefore a property of the problem, not of our proof technique.
@@
When $n\eta^{H-t}\Delta^{2}$ is small, $(1+\eta^{H-t}\Delta^{2})^{n}\approx 1+n\eta^{H-t}\Delta^{2}$, so the minimax error is close to $1/2$: the best possible test is barely better than a coin flip. For reinforcement learning, this means that critics trained from outcome labels cannot reliably evaluate steps beyond the critical horizon.
A natural question is whether parallel sampling, generating multiple independent trajectories from the same starting state, can overcome the exponential barrier. The next section shows that it cannot: width provides variance reduction but cannot amplify a signal that has already decayed.
@@ 4 Limits of Parallel Sampling
The Signal Decay Bound reveals that outcome signals weaken exponentially with depth. A natural response is to generate more samples: $W$ independent observations from the same starting condition, as in replicated testing or width-based methods like GRPO (DeepSeek-AI 2024). The intuition is that even if individual observations carry weak signal, averaging enough of them should recover the truth. This approach appears to extract more signal from the outcome data that Section 3 proved is fundamentally limited.
But does width truly overcome the exponential barrier, or merely work around it? This section shows that it does not: width reduces noise around an estimate but cannot amplify a signal that has already decayed. We proceed in three steps: derive variance reduction bounds under independence (Section 4.1), show that correlation limits effective width (Section 4.2), and synthesize practical guidance (Section 4.3). Proofs of all results in this section appear in Appendix 11.
@@ 4.1 Multi-Rollout Value Estimation
    Concentration (Hoeffding): With probability at least $1-\delta$:
    $|\hat{V}_{W}(s)-V^{\pi}(s)|\leq\sqrt{\frac{\ln(2/\delta)}{2W}}.$ (20)
@@
Example 4.3 (Width Requirements for Accurate Estimation)
To illustrate the concentration bound, suppose we want to estimate $V^{\pi}(s)$ to within $\pm 0.05$ with 95% confidence. Setting $\delta=0.05$ in Part 3 (so the bound holds with probability $1-\delta=0.95$) gives $\ln(2/\delta)=\ln(40)$. Requiring the error bound to be at most $0.05$:
$\sqrt{\frac{\ln(40)}{2W}}\leq 0.05\quad\Longrightarrow\quad W\geq\frac{\ln(40)}{2\times 0.0025}\approx 738.$ (21)
Approximately 740 rollouts per state suffice. This is a worst-case bound; when $V^{\pi}(s)$ is near 0 or 1, variance is smaller and fewer rollouts suffice. Relaxing to $\pm 0.1$ accuracy reduces the requirement to about 185 rollouts.
@@ 4.2 Correlation Limits Effective Width
To quantify this limitation, we must account for correlation among rollouts from the same prefix. Consider collecting data from $n$ independent starting points, generating $W$ rollouts from each. Rollouts from different starting points are independent, and their outcomes are uncorrelated. However, rollouts from the same starting point share the same prefix and tend to make similar errors, inducing positive correlation among their outcomes. Correlation arises from three principal mechanisms:
  1. 1.
    Shared prefix errors: When a prefix contains a latent mistake, all continuations inherit the defect. If an early assembly stage installs a component off-center, every subsequent inspection reveals the same misalignment; if a language model makes an algebraic error in step 3, all continuations from that point inherit it.
  2. 2.
    Mode collapse: Low-entropy policies generate nearly identical continuations from the same prefix ($\rho\to 1$). Low-temperature sampling may produce only 2–3 distinct reasoning paths from $W=50$ rollouts; a deterministic assembly line yields essentially identical outputs from the same inputs.
  3. 3.
    Systematic biases: Learned patterns produce consistent error modes activated by specific inputs. A model that consistently confuses correlation with causation will make the same inferential error across all rollouts from a given prefix, regardless of token-level variation.
When $W=1$ (one rollout per starting point), we obtain $n$ i.i.d. outcome observations. When $W>1$, this within-starting-point correlation reduces the effective sample size, as the following theorem quantifies.
Theorem 4.4 (Effective Width Under Correlation)
@@
$\mathrm{Var}(\hat{V}_{W}(s)\mid s)=\frac{\sigma^{2}(s)}{W}\cdot(1+(W-1)\rho),$ (22)
@@
Theorem 4.4 shows that correlation reduces effective width from $W$ to $W_{\mathrm{eff}}$. How does this affect the critical horizon? While the $W$ rollouts within each starting point are correlated, the $n$ starting points are independent. Define the effective sample size as $N_{\mathrm{eff}}:=n\cdot W_{\mathrm{eff}}$. Combined with the Signal Decay Bound (Theorem 3.9), distinguishing hypotheses about step $t$ requires
$N_{\mathrm{eff}}\;\gtrsim\;\frac{1}{\eta^{H-t}\Delta^{2}}.$ (24)
@@
Let $n$ be the number of independent starting points and $W_{\mathrm{eff}}$ the effective width per starting point. The critical horizon extends to
$H_{\mathrm{crit}}=\frac{\ln(n\cdot W_{\mathrm{eff}}\cdot\Delta^{2})}{\ln(1/\eta)}.$
@@
4.3 Summary: When Width Works
The preceding analysis yields two necessary conditions for width to be beneficial:
  1. 1.
    Signal must exist: The step must lie within the critical horizon ($H-t\leq H_{\mathrm{crit}}$). Beyond this, outcome data is uninformative regardless of sample size.
  2. 2.
    Rollouts must be approximately independent: Effective width saturates at $1/\rho$, so high correlation eliminates the benefit of additional rollouts.
When either condition fails, width provides diminishing returns. The following decision rule provides a practical roadmap for determining when width-based methods can succeed:
Practical Decision Rule for Width-Based Methods 1. Estimate correlation $\rho$ from pilot data or domain knowledge (e.g., deterministic policies have $\rho\approx 1$; high-temperature sampling has lower $\rho$). Compute $W_{\mathrm{eff}}=W/(1+(W-1)\rho)$. 2. Compute effective horizon: $H_{\mathrm{crit}}=\ln(n\cdot W_{\mathrm{eff}}\cdot\Delta^{2})/\ln(1/\eta)$, where $\Delta^{2}$ is the initial $\chi^{2}$-divergence between hypotheses and $\eta$ is the contraction coefficient. 3. Decision: If $H\leq H_{\mathrm{crit}}$, width-based methods can succeed. Otherwise, process supervision is necessary.
The next section addresses how step-level supervision, once obtained, should be aggregated into training objectives.
@@ 5.2 Gradient Decay in the Multiplicative Objective
For $\alpha\in[0,1]$, the interpolated objective is:
$J_{\alpha}(\theta):=(1-\alpha)\cdot J_{\mathrm{add}}(\theta)+\alpha\cdot J_{\mathrm{mult}}(\theta).$ (25)
At $\alpha=0$, the objective is learnable but incorrect; at $\alpha=1$, correct but potentially unlearnable. The following result characterizes how $\alpha$ should evolve with policy quality.
@@
Assume $\nabla_{\theta}J_{\mathrm{add}}$ and $\nabla_{\theta}J_{\mathrm{mult}}$ have non-negative inner product. To maintain $\|\nabla_{\theta}J_{\alpha}\|\geq c\cdot\|\nabla_{\theta}J_{\mathrm{add}}\|$ for some $c\in(0,1)$, it suffices to set $\alpha\leq 1-c$.
The proposition establishes that the interpolated objective preserves gradient signal when $\alpha$ is sufficiently small. In the independence model where $J_{\mathrm{add}}=Hp$ and $J_{\mathrm{mult}}=p^{H}$, the gradients satisfy $\|\nabla_{\theta}J_{\mathrm{mult}}\|=p^{H-1}\|\nabla_{\theta}J_{\mathrm{add}}\|$. As the policy improves and per-step success probability $p$ approaches $1$, this attenuation factor approaches $1$, allowing $\alpha$ to increase toward $1$ without sacrificing gradient magnitude. This yields a curriculum: begin training with $\alpha$ near 0 to ensure learnability, then gradually increase $\alpha$ toward $1$ as the policy improves, transitioning from a learnable but misspecified objective to one that is correctly specified.

### v2

@@ 3.3 The Signal Decay Bound
In applications: $H_{0}$ corresponds to “step $t$ is valid” (the transition $K_{t-1}^{(0)}$ produces correct states), while $H_{1}$ corresponds to “step $t$ contains an error” (the transition $K_{t-1}^{(1)}$ may produce erroneous states). In manufacturing: $H_{0}$ means stage $t$ operated correctly; $H_{1}$ means a defect was introduced at stage $t$. The shared-suffix condition isolates step $t$ as the sole source of difference: downstream stages process whatever they receive in the same way under both hypotheses, so any difference in the outcome distribution is attributable entirely to step $t$. Errors introduced at step $t$ may be corrected or may cascade through subsequent stages, but the stages themselves operate identically. Appendix 15 discusses how $\Delta^{2}$ can be estimated from pilot audits in operations settings. We index step $t$ as the abstract state $Z_{t}$ immediately after the $t$-th transition, so HtH-t counts the number of transitions remaining to reach the outcome.
@@
$\mathrm{err}^{*}(P^{(0)},P^{(1)}):=\inf_{\psi}\left(\mathbb{P}_{P^{(0)}}(\psi=1)+\mathbb{P}_{P^{(1)}}(\psi=0)\right)=1-d_{\mathrm{TV}}(P^{(0)},P^{(1)}),$}
where the infimum is over all measurable tests $\psi:\mathcal{Y}\to\{0,1\}$. For $n$ i.i.d. samples, the $n$-sample total testing error is $\mathrm{err}^{*}_{n}:=1-d_{\mathrm{TV}}(P^{(0)\otimes n},P^{(1)\otimes n})$, where the test acts on $\mathcal{Y}^{n}$. The hypotheses are $(n,\epsilon)$-distinguishable if $\mathrm{err}^{*}_{n}\leq\epsilon$.
Thus $(n,\epsilon)$-distinguishability requires $d_{\mathrm{TV}}(P^{(0)\otimes n},P^{(1)\otimes n})\geq 1-\epsilon$.
@@
    Sample complexity lower bound: For any target error $\epsilon\in(0,1/2)$, in the regime $\eta^{H-t}\Delta^{2}\leq 1$, any test achieving total testing error (Definition 3.7) at most $\epsilon$ requires
    $n\geq\frac{(1-\epsilon)^{2}}{\eta^{H-t}\cdot\Delta^{2}}$}
    i.i.d. outcome samples. When $\eta^{H-t}\Delta^{2}>1$, this lower bound becomes $O(1)$; in this regime, the outcome distributions are well-separated and $O(1)$ samples suffice to achieve any constant total error rate.
@@
    $n=\Omega\left(\left(\frac{1}{\eta}\right)^{H-t}\right).$}
Remark 11.4 in Appendix 11 extends Theorem 3.8 to noisy terminal observations $R\sim G(\cdot\mid Z_{H})$.
@@
For initial divergence $\Delta^{2}$, sample budget $n$, and any target error $\epsilon\in(0,1/2)$, a necessary condition for achieving total testing error at most $\epsilon$ is that the step lies within the $\epsilon$-critical horizon:
$H-t\leq H_{\mathrm{crit}}(\epsilon):=\max\!\left\{0,\;\frac{\ln\!\bigl(n\Delta^{2}/(1-\epsilon)^{2}\bigr)}{\ln(1/\eta)}\right\}.$}
Steps more than $H_{\mathrm{crit}}(\epsilon)$ from the outcome require exponentially many samples to evaluate from outcome labels alone. When $n\Delta^{2}<(1-\epsilon)^{2}$, the horizon is zero: even the final step cannot be distinguished at error $\epsilon$. For intuition, $H_{\mathrm{crit}}\approx\ln(n\Delta^{2})/\ln(1/\eta)$ when $\epsilon$ is a fixed constant.
@@
With $\eta=0.9$, $\Delta^{2}=0.1$, and $n=10^{6}$, the critical horizon is $H_{\mathrm{crit}}=\ln(10^{5})/\ln(1/0.9)\approx 110$: one million outcome samples can reliably evaluate steps up to 110 steps before the outcome. With stronger contraction $\eta=0.7$, the same budget yields $H_{\mathrm{crit}}\approx 32$. The formula depends logarithmically on $n$ and $\Delta^{2}$ but much more strongly on $\eta$: doubling $n$ extends the horizon by $\ln 2/\ln(1/\eta)$ steps ($\approx 6.6$ when $\eta=0.9$), while reducing $\eta$ from $0.9$ to $0.7$ cuts it by a factor of three.
The exponential rate in Theorem 3.8 is achievable: the following proposition constructs a Bernoulli endpoint-testing instance in which the likelihood ratio test matches the lower bound up to constants.
@@
Assume $p_{0}\in[\epsilon_{0},1-\epsilon_{0}]$ for some constant $\epsilon_{0}>0$, so the outcome is non-degenerate under both hypotheses. Then the $\chi^{2}$-divergence between outcome distributions is $\chi^{2}(P^{(0)}_{R}\|P^{(1)}_{R})=\Theta(\eta^{H-t}\Delta^{2})$, and the standard likelihood ratio test achieves total testing error $\epsilon$ with sample complexity $n=O(1/(\eta^{H-t}\Delta^{2}))$.
Together, Theorem 3.8 and Proposition 3.11 show that the exponential rate $(1/\eta)^{H-t}$ is both necessary and sufficient for the matched endpoint-testing problem: the scaling is tight up to constants. Example 11.8 in the appendix constructs a full Markov-chain instance achieving this rate with equality, confirming that the barrier is a property of the information structure, not of our proof technique.
@@
Corollary 3.12 uses the per-hypothesis error convention $\max_{i}\mathbb{P}_{i}(\psi\neq i)$, standard for Le Cam bounds, whereas Definition 3.7 uses the total error $\mathbb{P}_{0}(\psi=1)+\mathbb{P}_{1}(\psi=0)$; the two are related by a factor of two and yield equivalent scaling in $\eta^{H-t}\Delta^{2}$. When $n\eta^{H-t}\Delta^{2}$ is small, $(1+\eta^{H-t}\Delta^{2})^{n}\approx 1+n\eta^{H-t}\Delta^{2}$, so the minimax error is close to $1/2$: the best possible test is barely better than a coin flip. For reinforcement learning, this means that critics trained from outcome labels cannot reliably evaluate steps beyond the critical horizon.
@@ 4 Limits of Parallel Sampling
The Signal Decay Bound reveals that outcome signals weaken exponentially with depth. A natural response is to generate more samples: $W$ independent observations from the same starting condition, as in replicated testing or width-based methods like GRPO (DeepSeek-AI 2024). The intuition is that even if individual observations carry weak signal, averaging enough of them should recover the truth. This approach might seem to extract more signal from outcome data, despite Section 3 showing fundamental limits.
But can width truly overcome the exponential barrier? This section shows that it cannot: width reduces noise around an estimate but cannot amplify a signal that has already decayed. The statistical tools, variance reduction under averaging and design effects under equicorrelation (kish1965survey), are classical, but their interaction with the signal decay structure of Section 3 yields new limits on the reach of parallel sampling. We proceed in two steps: derive variance reduction bounds under independence (Section 4.1) and show that correlation limits effective width (Section 4.2). Proofs of all results in this section appear in Appendix 12.
@@ 4.1 Multi-Rollout Value Estimation
    Concentration (Hoeffding): With probability at least $1-\delta$: $|\hat{V}_{W}(s)-V^{\pi}(s)|\leq\sqrt{\frac{\ln(2/\delta)}{2W}}$
@@ 4.2 Correlation Limits Effective Width
To quantify this limitation, we must account for correlation among rollouts from the same prefix. Consider collecting data from $n$ independent starting points, generating $W$ rollouts from each. Rollouts from different starting points are independent, and their outcomes are uncorrelated. However, rollouts from the same starting point share the same prefix and tend to make similar errors, inducing positive correlation among their outcomes. Correlation arises from shared prefix errors (all continuations inherit a latent early defect), mode collapse (low-entropy policies generate near-identical continuations, driving $\rho\to 1$), and systematic biases (learned error patterns activated by specific inputs). When $W=1$ (one rollout per starting point), we obtain $n$ i.i.d. outcome observations. When $W>1$, this within-starting-point correlation reduces the effective sample size, as the following theorem quantifies.
Theorem 4.3 (Effective Width Under Correlation (cf. kish1965survey))
@@
$\mathrm{Var}(\hat{V}_{W}(s)\mid s)=\frac{\sigma^{2}(s)}{W}\cdot(1+(W-1)\rho),$}
@@
The equicorrelation variance formula in Theorem 4.3 is a classical design-effect result in survey sampling (kish1965survey); what is new is its implication for the credit assignment horizon, which we develop next. While the $W$ rollouts within each starting point are correlated, the $n$ starting points are independent. Define the effective sample size as $N_{\mathrm{eff}}:=n\cdot W_{\mathrm{eff}}$. Combined with the Signal Decay Bound (Theorem 3.8), distinguishing hypotheses about step $t$ requires
$N_{\mathrm{eff}}\;\gtrsim\;\frac{1}{\eta^{H-t}\Delta^{2}}.$}
@@
Let $n$ be the number of independent starting points and $W_{\mathrm{eff}}$ the effective width per starting point. For fixed $\epsilon$, the critical horizon extends to
$H_{\mathrm{crit}}=\frac{\ln(n\cdot W_{\mathrm{eff}}\cdot\Delta^{2})}{\ln(1/\eta)},$
where the $O(\ln(1/(1-\epsilon)))$ correction from Corollary 3.9 is suppressed.
@@
In sum, width helps only when the step lies within the critical horizon and rollouts are approximately independent; beyond either limit, additional rollouts yield diminishing returns. For deep processes where $H>H_{\mathrm{crit}}$, width cannot substitute for step-level supervision; Appendix 16 provides a decision rule: width alone is provably insufficient whenever $H>\ln(n\Delta^{2}/\rho)/\ln(1/\eta)$, and in practice one inspection dominates doubling $W$ whenever the maximal gap exceeds $H_{\mathrm{crit}}(\epsilon)$. This raises the question of how step-level supervision should be aggregated.
@@ 5.2 Gradient Decay in the Multiplicative Objective
For $\lambda\in[0,1]$, the interpolated objective is:
$J_{\lambda}(\theta):=(1-\lambda)\cdot J_{\mathrm{add}}(\theta)+\lambda\cdot J_{\mathrm{mult}}(\theta).$}
At $\lambda=0$, the objective is learnable but incorrect; at $\lambda=1$, correct but potentially unlearnable. The following result characterizes how $\lambda$ should evolve with policy quality.
@@
Assume $\nabla_{\theta}J_{\mathrm{add}}$ and $\nabla_{\theta}J_{\mathrm{mult}}$ have non-negative inner product. To maintain $\|\nabla_{\theta}J_{\lambda}\|\geq c\cdot\|\nabla_{\theta}J_{\mathrm{add}}\|$ for some $c\in(0,1)$, it suffices to set $\lambda\leq 1-c$.
The proposition establishes that the interpolated objective preserves gradient signal when $\lambda$ is sufficiently small. In the independence model where $J_{\mathrm{add}}=Hp$ and $J_{\mathrm{mult}}=p^{H}$, the gradients satisfy $\|\nabla_{\theta}J_{\mathrm{mult}}\|=p^{H-1}\|\nabla_{\theta}J_{\mathrm{add}}\|$. As the policy improves and per-step success probability $p$ approaches $1$, this attenuation factor approaches $1$, allowing $\lambda$ to increase toward $1$ without sacrificing gradient magnitude. This yields a curriculum: begin training with $\lambda$ near 0 to ensure learnability, then gradually increase $\lambda$ toward $1$ as the policy improves, transitioning from a learnable but misspecified objective to one that is correctly specified.